OpenStudy (anonymous):

I'm new to this site.......hope this is right way to write this. perform operation.... 3/x+2/3x^2

OpenStudy (shayaan_mustafa):

Hi Isteele :) How are you ? Welcome to Openstudy :D

OpenStudy (shayaan_mustafa):

Yes it is right my friend.

OpenStudy (anonymous):

it's not like i wanted it to look.

OpenStudy (shayaan_mustafa):

\[\large \frac{3}{x}+\frac{2}{3x^2}\] Is this your question look like?

OpenStudy (anonymous):

i am having a hard time with this class...i haven't had algebra, and struggling

OpenStudy (anonymous):

yes it is. could you help me with the steps to figuring the answer?

OpenStudy (shayaan_mustafa):

Don't worry my friend. We are here to help you dear.

OpenStudy (anonymous):

whoo hoo!!!!!

OpenStudy (shayaan_mustafa):

\[\large \frac{x^2+2}{3x^2}\] It is solved now. It is your answer.

OpenStudy (precal):